OpenStudy (anonymous):

What is the solution to the system of equations x=3y+5 x-4y=3

OpenStudy (anonymous):

1) x=3y +5 subtract 3y from both sides and you will get -3y-x=5 then subtract x from both sides and you will get -3y=-x+5 divide the whole equation by negative three your answer is \[y=-\frac{ 1 }{ 3 }x+5\] 2) x-4y=3 subtract x from both sides and you will get -4y=-x+3 divide the whole equation by negative four your answer is \[y=-\frac{ 1 }{ 4 }x+3\]

OpenStudy (anonymous):

the answer is an ordered pair though

OpenStudy (anonymous):

ohhh ok just a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

ya is it asking you to find either x or y?

OpenStudy (anonymous):

both

OpenStudy (anonymous):

ok so first you need to get the first equation switched around so it equals a number like the second equation x=3y+5 subtract 3y from both sides 3y+x=5 x-4y=3 (your second equation) then you need to first eliminate x by subtracting the two equations

OpenStudy (anonymous):

ok so you need to subtract them

OpenStudy (anonymous):

how do i do that @Sticky

OpenStudy (anonymous):

\[\frac{ x-3y=5 }{ x-4y=3 }=7y=2\]

OpenStudy (anonymous):

so y=7

OpenStudy (anonymous):

and x=2

OpenStudy (anonymous):

no \[y=\frac{ 2 }{ 7 }\]

OpenStudy (anonymous):

you had to divide 7y by 7 and 2 divieded by 7 to get y

OpenStudy (anonymous):

so y= 3.5

OpenStudy (anonymous):

yes now substitute 3.5 as Y in one of the equations to find x

OpenStudy (anonymous):

so x=22.5

OpenStudy (anonymous):

x=15.5

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

@Sticky

OpenStudy (anonymous):

I don't think so :/ I'm pretty sure I messed up somewhere you solved it right but there should only be one x just a minute

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@Gondola

OpenStudy (anonymous):

I think its (-5.5,3.5)

OpenStudy (anonymous):

how

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

http://www.purplemath.com/modules/systlin5.htm if its wrong you can find out how to do it here sorry for all the trouble :/

OpenStudy (anonymous):

thank you so much it help

OpenStudy (anonymous):

no probs cya